OpenStudy (anonymous):

Which values of p and q would make the value of the following expression equal to 29i? (2 – 5i)(p + q)i A.p = 2, q = 5i B.p = –2, q = 5i C.p = 2, q = –5i D.p = –2, q = –5i

OpenStudy (solomonzelman):

expand first: (2 – 5i)(p + q)i =

OpenStudy (anonymous):

Okay @SolomonZelman could probably explain exactly how to do this because he is smart like that but I was able to pass the SAT by plugging these things in so lets start with A.

OpenStudy (anonymous):

(2-5i)(2+5i)i 4 - 10i + 10i - 25i^2 4 - 10i + 10i - 25(-1) 4 - 0i - (-25) 4 + 25 29

OpenStudy (anonymous):

@SolomonZelman just check my problem on A so I know I am not giving @dreamstoogood wrong info.

OpenStudy (solomonzelman):

u need it to equal 29i, not 29

OpenStudy (anonymous):

o.o What did I do wrong?

OpenStudy (solomonzelman):

(2 – 5i)(p + q)i=29i (in a+bi form, it is 0+29i) (2 – 5i)(p + q)i = expands to >> 2p+2q-5pi-5qi so, 2p+2q-5pi-5qi=29i in a+bi form, (2p+2q)+(-5p-5q)i=0+29i therefore 2p+2q=0 (-5p-5q)i=29i

OpenStudy (solomonzelman):

so there is a system of equations to solve

OpenStudy (solomonzelman):

you can cancel i' in equation 2, and from that point and on it is just algebra.